OpenStudy (anonymous):

a block of mass M is pulled along a horizontal frictionless surface by a rope of mass m. if a force F is applied at one end of the rope, what is the force which the rope exerts on the block?

OpenStudy (yash2651995):

F.. action reaction pair.. (rope is taut, assumed.. jo ho gaa hi..)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

I think it will be a little less than F (depending upon mass of rope). This is because the tension in the rope is not constant as the rope is having mass and is accelerated under the force F and the force that the block applies on it (Let it be F1) Acceleration of system is: F/(M+m) So net force on rope = m*(F/(M+m)) {m*a} i.e. F-F1=m*(F/M+m) F1 =M*(F/M+m) And this is the required answer